Gandolfi Construction Co. purchased a CAT 336DL earth mover at a cost of $1,000,000 in January 2019. The companys estimatedb. Calculate what Gandolfis ROI would have been for the year ended December 31, 2019, had the company used the double- decli
Answer #1

(a)

Return on Investment for the year ended December 31, 2019

Choose Numerator

/

Choose Denominator

=

ROI

Net Income

/

Average Total Assets

=

ROI

$960000

/

$8000000

=

12%

(b)

Return on Investment for the year ended December 31, 2019

Choose Numerator

/

Choose Denominator

=

ROI

Net Income

/

Average Total Assets

=

ROI

$810000

/

$7925000

=

10.2%

Net Income = $960000 + Depreciation under SLM – Depreciation under Double declining balance

Depreciation under SLM = ($1000000 - $200000)/8years = $100000

Depreciation under Double declining balance:-

Rate = (1/8 years)* 2 = 25%

$1000000 * 25% = $250000

Net Income = $960000 + $100000 – $250000 = $810000

Average Total Asset = $8000000 – [($250000 - $100000)/2] = $7925000